IN 421 Investments  - Chapter 18 Homework

Questions:

6. You expect the price of a stock to decline but do not want to sell the stock short and run the risk that the price of the stock may rise dramatically. How could you use a bear spread strategy to take advantage of your expectation of a lower stock price?

7. You sell a stock short. How can you use an option to reduce your risk of loss should the price of the stock rise?

9. If you thought a stock was fairly valued and its price would not change, how could you use a straddle to take advantage of your valuation? If you follow this strategy and the stock’s price does not remain stable, have you increased your risk exposure?
Answer #1

6.

A bear spread is a strategy used in options trading in which a trader purchases a put option contract with a higher strike price and sells a put option contract with a lower strike price.

This strategy is used to maximize profit in case of a decline in stock price while still limiting any loss that could occur from a steep increase in price.

Let us take an example :-

Suppose current price of Stock A is $100

Let us assume that price of Put Option with a strike price of 100 is $10 and price of Put Option with a strike price of 90 is $5

If the trader thinks that price of Stock A will decline, he will enter into a bear spread by buying Put Option with a strike price of 100 for $10 and selling Put Option with a strike price of 90 for $5

The strategy will result into an initial outflow of $(10-5) = $5

The maximum potential payoff during the time of expiry from this strategy will be the difference between 2 strike prices = (100-90) = $10

Hence, maximum possible gain from this strategy = $(10 - 5) = $5

Maximum possible loss in case of significant increase in the stock price will be the initial outflow of $5. Hence, this strategy will protect the trader from the risk of dramatic rise in stock price